OpenStudy (anonymous):

sin^2x-sinx= cos^2x solve for x a)over the domain of 0_

myininaya (myininaya):

\[\sin^2x-sinx=1-\sin^2x\] \[2\sin^2x-sinx-1=0\] \[u=sinx\] \[2u^2-u-1=0\] \[2u^2-2u+u-1=0\] \[2u(u-1)+(u-1)=0\] \[(u-1)(2u+1)=0\] \[u=sinx=1\] \[u=sinx=\frac{-1}{2}\]

OpenStudy (anonymous):

too fast o.O